OpenStudy (anonymous):

PLEASE HELP ME PLEASE PROBABILITY DISTRIBUTION AND EXPECTED VALUE QUESTION An online gambling site offers a first prize of $50,000 and two second prizes of $10,000 each for registered users when they place a bet. A random bet will be selected over a 24-hour period. One million bets are received in the contest. Find the expected winnings if you can place one registered bet of $1 in the given period. @amistre64

OpenStudy (kropot72):

The average value of the prizes is (50000 + 10000 + 10000)/3 = $23,333.33 The probability of a loss is (10^6 - 1)/10^6. So the expected loss is \[-1\times\frac{10^{6}-1}{10^{6}}\] The probability of a win is 1/10^6. So the expected gain is \[23333.33\times\frac{1}{10^{6}}\] The expected winnings is given by \[\frac{23333.33}{10^{6}}-\frac{10^{6}-1}{10^{6}}=you\ can\ calculate\] Hint: The expected winnings will be a loss!

OpenStudy (anonymous):

PLEASE WAT DOES THE 10^6 MEAN. AND HOW DID YOU GET IT ..@kropot72

OpenStudy (anonymous):

secondly how did you get the additional 10,000 in the begiinig of your work? @kropot72

OpenStudy (kropot72):

\[1,000,000=10^{6}\] One million can also be written as 10^6 in OpenStudy's convention. There are three prizes, one prize of $50,000 and two prizes of $10,000. the total prize money is 50,000 + 10,000 + 10,000 = $70,000. $70,000 is divided by three to get the average prize.

OpenStudy (anonymous):

please how did u get -1 in the expected loss calculation? @kropot72

OpenStudy (kropot72):

The amount of the bet is $1. The $1 is kept by the gambling site, so it is effectively a loss and is given a negative sign.

OpenStudy (anonymous):

oh ok,so wat i get from your last calculation should be my answer.?

OpenStudy (kropot72):

What answer do you get when you do the calculation?

OpenStudy (anonymous):

0.0233 @kropot72

OpenStudy (anonymous):

is it right? @kropot72

OpenStudy (kropot72):

The value of the first quotient is \[\frac{23333.33}{10^{6}}=0.02333333\] Now you need to subtract the value of the second quotient: \[0.02333333-\frac{999,999}{1,000,000}=you\ can\ calculate\] This will give a negative answer as a decimal fraction of a dollar.

OpenStudy (anonymous):

-0.9767

OpenStudy (anonymous):

@kropot72

OpenStudy (kropot72):

Good work! My result was an expected loss of 97.666 cents.

OpenStudy (anonymous):

u had no negative ? or i can also choose to write it as how you have written it,..and the question which says find the expected winnings doesnot necessarily mean a profit right??

OpenStudy (anonymous):

@kropot72

OpenStudy (kropot72):

Sorry for confusing you by expressing the answer as an expected loss. The question asks for the expected winnings, therefore I suggest that you write the answer as either -97.666 cents or as -$0.97666. Your choice might depend on the format provided for your answer.

OpenStudy (kropot72):

@payin Please ask if you need more help.

OpenStudy (anonymous):

oh ok.i would need to get over the solution a step at a time.

OpenStudy (kropot72):

I think that my first posting and the replies to your questions show all the steps. Which steps do you want me to go over?

OpenStudy (anonymous):

ive got it perfectly now..thanks

OpenStudy (kropot72):

You're welcome :)